OpenStudy (idealist10):

How to integrate (z+1)/(z-1)^2 dz?

OpenStudy (idealist10):

@ParthKohli @Abhisar @Destinymasha

OpenStudy (anonymous):

partial fractions

OpenStudy (idealist10):

@dumbcow

OpenStudy (dumbcow):

lets see i think a simple substitution will work here as well u = z - 1 du = dz --> z+1= (u+1)+1 = u+2 \[\int\limits \frac{u+2}{u^2} du = \int\limits \frac{1}{u} + 2 u^{-2}\]

OpenStudy (idealist10):

Thank you, @dumbcow